a rifle can shoot a 4.20g bullet at a speed of 965m/s

what is the kinetic energy of the bullet


1
Expert's answer
2021-05-30T13:31:41-0400

By definition, the kinetic energy is:


"K = \dfrac{mv^2}{2}"

where "m = 4.20g = 4.2\times 10^{-3}kg" is the mass of the bullet, and "v = 965 m\/s" is its speed. Thus, obtain:


"K = \dfrac{4.2\times 10^{-3}\cdot 965^2}{2} \approx 1.96\times 10^{6}J"

Answer. "1.96\times 10^{6}J".